Fill in the blanks.
 
(c) 12 + (–12) = ____
 
 
iaemfuewgnuersiojmhnbiorjhmneru

Grade:12th pass

3 Answers

Harshit Singh
2 years ago
Dear Student



(c) 12 + (-12) = 0



​Thanks
Pooventhra A
2 years ago
12+ ( -12 ) =0 Because, solve the bracket there is a plus sign outer the bracket and inside the bracket minus 12 is there.so solving the sign plus mutiply by minus sign results minus 12.now 12-12=0.
Sailikith Nallabirudu
2 years ago
12+(-12) is zero because ,if there is a negative number and postive number we subtract so 12+(-12) is zero